OpenStudy (anonymous):

Find the angle of elevation e: (3x+1) degree sign, d:2(x+8) degree sign

zepdrix (zepdrix):

So here I've drawn an angle of elevation, and an angle of depression. Hopefully the diagram makes sense.

zepdrix (zepdrix):

These angles turn out to be congruent since they are `Alternate Interior Angles`.

zepdrix (zepdrix):

Therefore: \(\Large\rm 2(x+8) =3x+1\) And solve for x! :)

OpenStudy (anonymous):

I got x =1 ???

zepdrix (zepdrix):

Hmmm that doesn't sound quite right :o

zepdrix (zepdrix):

\[\Large\rm 2(x+8) =3x+1\]Distributing the 2,\[\Large\rm 2x+16=3x+1\]

zepdrix (zepdrix):

How do we get the x's on the same side? What to do next? :o

OpenStudy (anonymous):

subtract 2x from both sides

OpenStudy (anonymous):

would x be 17?